# Exploring Multidimensional Checkworthiness: Designing AI-assisted Claim Prioritization for Human Fact-checkers

### v1

@@ 4.3. Targeted versus Abstract Prompting for Defining New Checkworthy Dimensions
RQ2.1 explores how participants created customized LLM filters, and the corresponding benefits and limitations. Customizable, user-generated filters enable users to filter for additional checkworthiness dimensions when keyword search and natively-supported checkworthy filters prove insufficient. When participants write LLM prompts in natural language, we can also explicitly understand what potential needs they have and how they express these needs.
Prompts written by participants ranged from specific to abstract, including: 1) targeted prompts to retrieve claims with specific topics or particular types of claims; and 2) abstract prompts as benchmark relevance criteria for claim exploration. We categorize these prompts according to different user intents (Table 5), which help us understand reasons participants created them. These intents also emphasize both specific and general fact-checking information needs given their familiarity of the fact-checking topics. We now proceed to describe how participants used LLMs to write targeted and abstract prompts.
@@
However, this also requires participants to write sufficient details in the prompt to ensure LLMs can retrieve targeted results accurately. Many participants (P4, 9, 10) noted that there is a learning curve to writing effective prompts (Giray, 2023), perhaps akin to effort required to write effective queries in early days of search engines. As search engines increasingly exploit LLM technology to better interpret user queries (goo, 2024), the added value of having a separate mechanism for writing LLM prompts as custom filters may correspondingly diminish. See further discussion in Section 5.
@@
Additionally, P1 created the prompt “chaos” to identify claims that broadly hint at potential harmful outcomes. This participant also created another prompt, “confusion,” aimed at uncovering “claims aimed at misleading the public about the effectiveness of prevention measures.
Participants said these abstract prompts could be very useful for exploratory search, especially when unfamiliar with specific news topics. Their LLM goal in such cases was to explore what potential claims LLMs could offer them based on high-level relevance factors, including the semantics or outcomes suggested by the claims, as a starting point for pursuing more specific, targeted claims:
I think that the LLM could be more useful if you’ve had less experience or you want to have a more broader idea than a specific query to [let it] find things for you.” (P12) “For more general things like news that emerged recently, if I don’t know what kind of information I’m searching for, giving a pointer instruction to the results could be very helpful.” (P15)
@@ 4.3.3. Comparing written prompts with existing checkworthy factors
However, a notable distinction in the written prompts is that participants would intentionally integrate the fact-checking topic, such as COVID-19, into the description of these checkworthy factors. For example, P9 created a customized facet called “Vaccine causes harm” (similar to the existing harmful filter) but elaborated: “Claims that vaccines are dangerous, often based on unreliable information, frequently include narratives that a specific group intends to cause harm or lead to death.” This highlights that while checkworthiness can be defined as several general relevant factors useful for exploratory claim searches, fact-checkers tend to contextualize these factors based on their familiarity with the topic, making them more topically relevant. Unlike traditional preset filters, which depend on curated datasets to train predictive models and may be less accurate when applied to out-of-distribution test datasets, participants can create LLM filters that incorporate richer contextual information, including insights that might be underrepresented in the training data.
Additionally, we also found that certain factors from Table 1, such as “Already checked,” “Amplification,” and “Difficulty,” were not explicitly mentioned by participants in the prompts. Since this analysis is a post-hoc comparison, we were unable to ask fact-checkers whether these factors were overlooked intentionally or for other reasons. One hypothesis could be that for “Already checked” claims, fact-checkers may have used the topical search or written targeted prompts to see if our dataset contains those already checked claims. In contrast, factors like “Amplification” and “Difficulty” might be perceived as too abstract. Fact-checkers might doubt the LLM ability to reliably capture these factors, underscoring the need for further investigation.
@@ 4.4.1. Usage behaviors in claims exploration
##\## Queries 2.00(2.48) 2.56(3.41) 10.0(0.256)
##\## Checkworthy slider changes 1.38(0.96) 6.44(5.84) 0.0(0.001)
@@
Table 7. Comparing behavioral measures (Section 3.4.4) with vs.  without customized filters for unidimensional and multidimensional interfaces. Mean (M) and standard deviation (SD) results are bolded if statistically significant for the Wilcoxon signed-rank test at p ¡ 0.05. Results show that the number of queries and query similarity slider changes significantly decrease in the multidimensional interface with the customized LLM slider. A decrease was also seen with the unidimensional interface but was not significant.
@@ 4.4.2. Success in finding checkworthy claims
To address potential inefficiency in selecting claims using multidimensional checkworthiness in the exploratory phase, future work might initially prioritize AI-recommended claims that meet different dimensions to reduce effort required for claim exploration (see Section 4.2). Once fact-checkers have quickly investigated these claims, they could then use the multi-faceted sliders to triage claims involving trade-offs between dimensions.
@@ 4.4.3. Subjective reflections
We analyzed participant self-reported metrics for their exploratory search experience between the unidimensional and multidimensional interface (Table 9). The multidimensional interface scored significantly higher in several aspects: “Understand topic scope” ($M_{uni}=3.69$, $M_{multi}=4.44$), “Search specific topic” ($M_{uni}=3.44$, $M_{multi}=4.56$), “Lookup many claims” ($M_{uni}=3.44$, $M_{multi}=4.25$), “Investigate multiple criteria” ($M_{uni}=3.38$, $M_{multi}=4.62$), and “Operationalize multiple criteria” ($M_{uni}=3.69$, $M_{multi}=4.12$). Overall, participants considered the multidimensional interface more useful in exploratory claim search (despite finding more unique claims with the unidimensional interface). The multidimensional interface also received a high score regarding the perceived usefulness and ease of use to support claim prioritization (see details in Appendix H).
@@ 5. Discussion
We begin by discussing design implications of fact-checker dynamic and hierarchical claim prioritization process (Section 5.1). Specifically, we summarize findings that reveal differences in fact-checker claim triage compared to prior research, highlighting underexplored aspects of their workflows (Section 5.1.1). Building on these insights, we then propose design suggestions for personalized and efficient tool supports based on fact-checker feedback (Section 5.1.2). Furthermore, we reflect on the dynamic nature of hierarchical relevance in claim prioritization and discuss its broader implications for other user activities (Section 5.1.3).
@@
Finally, we describe the study limitations identified during our RtD process. We hope this helps guide future researchers in exploring the multidimensional checkworthiness and developing more advanced claim prioritization tools (Section 5.3).
@@ 5.1.1. Summary of fact-checker claim prioritization process
As discussed in the related work (Section 2.2), several open questions remain in understanding fact-checker claim prioritization. For example, prior studies have yet to clarify the relative importance of different checkworthiness dimensions, and potential gaps between fact-checker self-reported perspectives and their actual behaviors involved in claim search and selection. By employing an RtD approach, which allows us to discover nuanced user practice and design knowledge grounded in making and doing, we have identified insights that further extend findings from earlier research.
@@
We also observed an important contrast with prior work. In particular, while fact-checkers might describe their claim prioritization as unstructured or less systematic (in the context of searching, browsing, and filtering claims across multi-dimensional checkworthiness), their interactions with our prototype actually reflected a hierarchical claim triage process used to enhance work efficiency (as Section 4.2). Moreover, by examining different intents reflected from their written LLM prompts, we found that fact-checker familiarity with fact-checking topics significantly influences the strategies they employ when using LLMs to create customized filters (described in Section 4.3).
@@ 5.1.2. Design implications for personalized and efficient claim triage
Streamlining user hierarchical claim prioritization. The hierarchical approach to claim prioritization reflects an underlying systematic process. Although less overt, this inherent structure can inform tool designs to help participants streamline their claim selection workflows. For example, if a claim mutually satisfies multiple checkworthy factors along with the prerequisite or important dimensions, this should be clearly shown in the interface. For example, P3 was interested in the “Likely false” and “Verifiable”. The tool should prioritize or separate claims that met both dimensions. P7 also remarked that some claims consistently ranked at the top across even when different dimensions were favored and suggested the tool could “display them at the very top end so it was easier for me to review.” Thus, beyond providing personalized weighting mechanism for users to prioritize their prerequisite or important dimensions, future designs could also allow fact-checkers to directly preset these priorities. The display of ranked claims should then explicitly indicate how each claim aligns with these priorities.
On the other hand, sometimes there may be a trade-off between competing checkworthiness dimensions. For example, P1 found that some claims are “flagged as very harmful but less interesting to the public.” Similar to P3 and P7 above, he suggested making such tradeoffs clearly apparent to allow him to validate whether AI predictions for those checkworthy dimensions matched human judgment. If there is a mismatch between human judgment and AI predictions, the weighting mechanism could become particularly valuable, enabling fact-checkers to reduce AI influence.
In prior work, Sehat et al. (2024) propose a conceptual framework for claim prioritization, which introduces a new and structured approach. However, its effective implementation heavily relies on fact-checker training and education. We expect our findings and design recommendations will complement this educational approach by specifying the AI tool supports needed to streamline claim triage where different checkworthiness dimensions are either satisfied or in competition with one another.
@@
First, when news event emerge and fact-checkers are uncertain on what topics or dis/misinformation narratives would be the central of fact-checking, they create general faceted filters (i.e., abstract prompts), which help understand the semantics or potential harmful outcomes of problematic information. These criteria are frequently-used filters across different news events, as P14 explained “[they are] pretty broad category often used to filter out [less important] claims.” As fact-checkers become more familiar with specific dis/misinformation narratives, direct LLM-integrated search would be more helpful as it can retrieve specific claims based on precise narratives (i.e., targeted prompts).
@@
As search queries and/or LLM prompts become more complex, it becomes more important to explain how search results relate to different portions of an input query/prompt (i.e., algorithmic transparency). In IR, this is typically the domain of query-biased summarization (Sarwar et al., 2021) or search result snippets that serve to explain how each result relates to user query. With our simple LLM integration, P2 reflected that “the result might not be directly explainable based on each of the things I wrote down [in the prompt]. ” P7 similarly remarked, “it would be nice to tell me why things were being arranged and returned in a certain way. For example, if I had looked up the spike protein in the prompt, I would have seen claims highlighting the spike protein.” While users today are very familiar with standard search engines and reformulating their queries to improve their search results, LLMs and prompt engineering are simultaneously far less familiar while also being seemingly far more complex.
This feedback highlights the importance of designing future LLM interfaces to help users better understand two key aspects: 1) how LLMs interpret lengthy prompts and demonstrate their alignment with user needs, and 2) how they retrieve, organize, or generate results to address these needs. Some existing technical research might help shed light on the interface design, such as work on decomposing complex claims to retrieve evidence for claim verification (Chen et al., 2022) and explaining how different parts of a prompt influence the LLM output (Feng et al., 2023). For example, if we could similarly decompose complex prompts into smaller portions and illustrate the salience of each portion on the LLM output, this could be very informative to users. This would not only enable users to refine their prompts more precisely but also help address potential transparency and trust issues with LLM-generated content.
@@ 5.1.3. The dynamic nature of hierarchical relevance in user information-seeking
The changing news environment influences how journalists dynamically evaluate what news is worth reporting. Our participants argued that assessing the relative importance of multidimensional checkworthiness — the hierarchy we identified from user patterns — would change with different news contexts because fact-checking coverage evolves as events unfold. P3 said, “between these factors, such as likely harm, spread, and topicality, if you ask me on different days, which of those I think is more important, I could give you a different answer.” He further explained that “for the last two months, we solely focus on checking claims around Israel and Gaza, [where] topic and public interest were more important [but] during the height of the COVID, the ability and likelihood to cause harm is the most prominent factor.” Checkworthy dimensions following the hierarchy identified from user patterns may thus change depending on world events, necessitating a flexible design mechanism for claim prioritization. More generally, priorities over different dimensions of checkworthiness (e.g., Table 1) can be expected naturally and dynamically vary over diverse contexts such as time, location, and by organizational and individual preferences. Beyond fact-checking, such dynamism also reflects a broader phenomenon familiar with other information seeking contexts that are more personal, subjective, and situational, and where the relative importance of particular search criteria changes and new criteria emerge (Schamber et al., 1990).
@@ 5.2.1. Claim prioritization as a shift in fact-checking focus
Over the past decade, changes in the media landscape, e.g., the surge of online social media, and news consumption behavior, e.g., the trend of reading news via social media platforms, have significantly transformed journalistic fact-checking. This transformation is also evident in the shift of claim prioritization from focusing solely on checking political claims to addressing online misinformation by considering a broader range of checkworthy factors (described in Section 2.1). Thus, claim prioritization has become a central challenge for fact-checkers. In this section, by illustrating the historical changes of claim prioritization and its difference across various fact-checking entities, we discuss how our study insights better inform the development of advanced tools in order to adapt to this evolving practice.
Traditionally, fact-checking was an ad-hoc process embedded within news production prior to publication, where editors worked closely with authors to verify the accuracy of arguments (Smith, 2003). Ideally, every claim in ready-to-publish news was considered important to fact-check as part of internal journalistic work. This helps maintain the credibility and reputation of the media outlet. As fact-checking has evolved into a post-hoc activity focused on assessing the accuracy of existing public statements, claim prioritization became a distinct step. In this step, journalists start assessing the newsworthiness of various claims and only select some of them that worth checking. While this step is recognized as important, it is not generally viewed as particularly challenging. For example, as highlighted in the ethnographic work of Graves (2016), post-hoc fact-checking was largely carried out by traditional newsrooms and media outlets that primarily focused on political claims. Journalists in these organizations often focus on statements made by politicians, other journalists, and public figures. This emphasis remains evident, as noted by some participants (P5, 12) and other related work (Mena, 2019; Micallef et al., 2022).
@@
Claim prioritization tool has the potential to accelerate the collaborative effort to meet fact-checking needs from different stakeholders. For example, by using the tool to prioritize claims based on organizational checkworthiness, fact-checkers across these organizations can enhance their capacity to address a wider range of claims. Additionally, the tool might help address existing ethical considerations, particularly when fact-checkers navigate competing objectives in selecting claims. According to our participants, these competing objectives often stem from their partnerships with social media platforms and regional differences. We discuss these in details in the following sections.

### v2

@@ 4.3. Targeted versus Abstract Prompting for Defining New Checkworthy Dimensions
RQ2.1 explores how participants created customized LLM filters and the corresponding benefits and limitations. Customizable, user-generated filters enable users to filter for additional checkworthiness dimensions when keyword search and natively-supported checkworthy filters prove insufficient. When participants write LLM prompts in natural language, we can also explicitly understand what potential needs they have and how they express these needs.
Prompts written by participants ranged from specific to abstract, including 1) targeted prompts to retrieve claims with specific topics or particular types of claims and 2) abstract prompts as benchmark relevance criteria for claim exploration. We categorize these prompts according to different user intents (Table 5), which helps us understand the reasons participants created them. These intents also emphasize both specific and general fact-checking information needs, given their familiarity with the fact-checking topics. We now proceed to describe how participants used LLMs to write targeted and abstract prompts.
@@
However, this also requires participants to write sufficient details in the prompt to ensure LLMs can retrieve targeted results accurately. Many participants (P4, 9, 10) noted that there is a learning curve to writing effective prompts (Giray, 2023), perhaps akin to the effort required to write effective queries in the early days of search engines. As search engines increasingly exploit LLM technology to better interpret user queries (goo, 2024), the added value of having a separate mechanism for writing LLM prompts as custom filters may correspondingly diminish. See further discussion in Section 5.
@@
Additionally, P1 created the prompt “chaos” to identify claims that broadly hint at potentially harmful outcomes. This participant also created another prompt, “confusion,” aimed at uncovering “claims aimed at misleading the public about the effectiveness of prevention measures.
Participants said these abstract prompts could be very useful for exploratory search, especially when they are unfamiliar with specific news topics. Their LLM goal in such cases was to explore what potential claims LLMs could offer them based on high-level relevance factors, including the semantics or outcomes suggested by the claims, as a starting point for pursuing more specific, targeted claims:
I think that the LLM could be more useful if you’ve had less experience or you want to have a broader idea than a specific query to [let it] find things for you.” (P12) “For more general things like news that emerged recently, if I don’t know what kind of information I’m searching for, giving a pointer instruction to the results could be very helpful.” (P15)
@@ 4.3.3. Comparing written prompts with existing checkworthy factors
However, a notable distinction in the written prompts is that participants would intentionally integrate the fact-checking topic, such as COVID-19, into the description of these checkworthy factors. For example, P9 created a customized facet called “Vaccine causes harm” (similar to the existing harmful filter) but elaborated on: “Claims that vaccines are dangerous, often based on unreliable information, frequently include narratives that a specific group intends to cause harm or lead to death.” This highlights that while checkworthiness can be defined as several general relevant factors useful for exploratory claim searches, fact-checkers tend to contextualize these factors based on their familiarity with the topic, making them more topically relevant. Unlike traditional preset filters, which depend on curated datasets to train predictive models and may be less accurate when applied to out-of-distribution datasets, participants can create LLM filters that incorporate richer contextual information, including insights that might be underrepresented in the training data.
Additionally, we also found that certain factors from Table 1, such as “Already checked,” “Amplification,” and “Difficulty,” were not explicitly written by participants in the prompts. Since this analysis is a post-hoc comparison, we were unable to ask fact-checkers whether these factors were overlooked intentionally or for other reasons. One hypothesis could be that for “Already checked” claims, fact-checkers have already used the topical search or written targeted prompts to see if the dataset contains those already checked claims. In contrast, factors like “Amplification” and “Difficulty” might be perceived as too abstract. Fact-checkers might doubt the LLM ability to reliably capture these factors, underscoring the need for further investigation.
@@ 4.4.1. Usage behaviors in claims exploration
##\## Queries 2.00(2.48) 2.56(3.41) 10.00(0.25)
##\## Checkworthy slider changes 1.38(0.96) 6.44(5.84) 0.00(0.00)
@@
Table 7. Comparing behavioral measures (Section 3.4.4) with vs.  without customized filters for unidimensional and multidimensional interfaces. Mean (M) and standard deviation (SD) results are bolded if statistically significant for the Wilcoxon signed-rank test at p ¡ 0.05. Results show that the number of queries and query similarity slider changes significantly decrease in the multidimensional interface with the customized LLM slider. A decrease was also seen with the unidimensional interface, but it was not significant.
@@ 4.4.2. Success in finding checkworthy claims
To address potential inefficiency in selecting claims using multidimensional checkworthiness in the exploratory phase, future work might initially prioritize AI-recommended claims that meet different dimensions to reduce the effort required for claim exploration (see Section 4.2). Once fact-checkers have quickly investigated these claims, they could then use the multi-faceted sliders to triage claims involving trade-offs between dimensions.
@@ 4.4.3. Subjective reflections
We analyzed participant self-reported metrics for their exploratory search experience between the unidimensional and multidimensional interface (Table 9). The multidimensional interface scored significantly higher in several aspects: “Understand topic scope” ($M_{uni}=3.69$, $M_{multi}=4.44$), “Search specific topic” ($M_{uni}=3.44$, $M_{multi}=4.56$), “Lookup many claims” ($M_{uni}=3.44$, $M_{multi}=4.25$), “Investigate multiple criteria” ($M_{uni}=3.38$, $M_{multi}=4.62$), and “Operationalize multiple criteria” ($M_{uni}=3.69$, $M_{multi}=4.12$). Overall, participants considered the multidimensional interface more useful in exploratory claim searches (despite finding more unique claims with the unidimensional interface). The multidimensional interface also received a high score regarding the perceived usefulness and ease of use to support claim prioritization (see details in Appendix H).
@@ 5. Discussion
We begin by discussing the design implications of the fact-checker dynamic and hierarchical claim prioritization process (Section 5.1). Specifically, we summarize findings that reveal differences in fact-checker claim triage compared to prior research, highlighting underexplored aspects of their workflows (Section 5.1.1). Building on these insights, we then propose design suggestions for personalized and efficient tool supports based on fact-checker feedback (Section 5.1.2). Furthermore, we reflect on the dynamic nature of hierarchical relevance in claim prioritization and discuss its broader implications for other user activities (Section 5.1.3).
@@
Finally, we describe the study limitations identified during our RtD process. We hope this helps guide future researchers in exploring multidimensional checkworthiness and developing more advanced claim prioritization tools (Section 5.3).
@@ 5.1.1. Summary of fact-checker claim prioritization process
As discussed in the related work (Section 2.2), several open questions remain in understanding fact-checker claim prioritization. For example, prior studies have yet to clarify the relative importance of different checkworthiness dimensions and potential gaps between fact-checker self-reported perspectives and their actual behaviors involved in claim search and selection. By employing an RtD approach, which allows us to discover nuanced user practice and design knowledge grounded in making and doing, we have identified insights that further extend findings from earlier research.
@@
We also observed an important contrast with prior work. In particular, while fact-checkers might describe their claim prioritization as unstructured or less systematic (in the context of searching, browsing, and filtering claims across multi-dimensional checkworthiness), their interactions with our prototype actually developed a hierarchical claim triage process to enhance work efficiency (in Section 4.2). Moreover, by examining different intents reflected from their written LLM prompts, we found that fact-checker familiarity with fact-checking topics significantly influenced the strategies they employed when using LLMs to create customized filters (described in Section 4.3).
@@ 5.1.2. Design implications for personalized and efficient claim triage
Streamlining user hierarchical claim prioritization. The hierarchical approach to claim prioritization reflects an underlying systematic process. Although less overt, this inherent structure can inform tool designs to help participants streamline their claim selection workflows. For example, if a claim mutually satisfies multiple checkworthy factors along with the prerequisite or important dimensions, this should be clearly shown in the interface. For example, P3 was interested in the “Likely false” and “Verifiable”. The tool should prioritize or separate claims that meet both dimensions. P7 also remarked that some claims consistently ranked at the top across even when different dimensions were favored and suggested the tool could “display them at the very top end so it was easier for me to review.” Thus, beyond providing a personalized weighting mechanism for users to prioritize their prerequisite or important dimensions, future designs could also allow fact-checkers to directly preset these priorities. The display of ranked claims should then explicitly indicate how each claim aligns with these priorities.
On the other hand, sometimes, there may be a trade-off between competing checkworthiness dimensions. For example, P1 found that some claims are “flagged as very harmful but less interesting to the public.” Similar to P3 and P7 above, he suggested making such trade-offs clearly apparent to allow him to validate whether AI predictions for those checkworthy dimensions matched human judgment. If there is a mismatch between human judgment and AI predictions, the weighting mechanism could become particularly valuable, enabling fact-checkers to reduce AI influence.
In prior work, Sehat et al. (2024) propose a conceptual framework for claim prioritization, which introduces a new and structured approach. However, its effective implementation heavily relies on fact-checker training and education. We expect our findings and design recommendations will complement this educational approach by specifying the AI tool supports needed to streamline claim triage, especially where different checkworthiness dimensions are either satisfied or in competition with one another.
@@
First, when news events emerge and fact-checkers are uncertain about what topics or dis/misinformation narratives would be the central of fact-checking, they create general faceted filters (i.e., abstract prompts) to understand the semantics or potential harmful outcomes of problematic information. These criteria are frequently-used filters across different news events, as P14 explained “[they are] pretty broad category often used to filter out [less important] claims.” As fact-checkers become more familiar with specific dis/misinformation narratives, direct LLM-integrated search would be more helpful as it can retrieve specific claims based on precise narratives (i.e., targeted prompts).
@@
As search queries and/or LLM prompts become more complex, it becomes more important to explain how search results relate to different portions of an input query/prompt (i.e., algorithmic transparency). In IR, this is typically the domain of query-biased summarization (Sarwar et al., 2021) or search result snippets that serve to explain how each result relates to a user query. With our simple LLM integration, P2 reflected that “the result might not be directly explainable based on each of the things I wrote down [in the prompt]. ” P7 similarly remarked, “it would be nice to tell me why things were being arranged and returned in a certain way. For example, if I had looked up the spike protein in the prompt, I would have seen claims highlighting the spike protein.” While users today are very familiar with standard search engines and reformulating their queries to improve their search results, LLMs and prompt engineering are simultaneously far less familiar while also being seemingly far more complex.
This feedback highlights the importance of designing future LLM interfaces to help users better understand two key aspects: 1) how LLMs interpret lengthy prompts and demonstrate their alignment with user needs and 2) how they retrieve, organize, or generate results to address these needs. Some existing technical research might help shed light on the interface design, such as work on decomposing complex claims to retrieve evidence for claim verification (Chen et al., 2022) and explaining how different parts of a prompt influence the LLM output (Feng et al., 2023). For example, if we could similarly decompose complex prompts into smaller portions and illustrate the salience of each portion on the LLM output, this could be very informative to users. This would not only enable users to refine their prompts more precisely but also help address potential transparency and trust issues with LLM-generated content.
@@ 5.1.3. The dynamic nature of hierarchical relevance in user information-seeking
The changing news environment influences how journalists dynamically evaluate what news is worth reporting. Our participants argued that assessing the relative importance of multidimensional checkworthiness — the hierarchy we identified from user patterns — would change with different news contexts because fact-checking coverage evolves as events unfold. P3 said, “between these factors, such as likely harm, spread, and topicality, if you ask me on different days, which of those I think is more important, I could give you a different answer.” He further explained that “for the last two months, we solely focus on checking claims around Israel and Gaza, [where] topic and public interest were more important [but] during the height of the COVID, the ability and likelihood to cause harm is the most prominent factor.” Checkworthy dimensions following the hierarchy identified from user patterns may thus change depending on world events, necessitating a flexible design mechanism for claim prioritization. More generally, priorities over different dimensions of checkworthiness (e.g., Table 1) can be expected to naturally and dynamically vary over diverse contexts such as time, location, and organizational and individual preferences. Beyond fact-checking, such dynamism also reflects a broader phenomenon familiar with other information-seeking contexts that are more personal, subjective, and situational, and where the relative importance of particular search criteria changes and new criteria emerge (Schamber et al., 1990).
@@ 5.2.1. Claim prioritization as a shift in fact-checking focus
Over the past decade, changes in the media landscape, e.g., the surge of online social media, and news consumption behavior, e.g., the trend of reading news via social media platforms, have significantly transformed journalistic fact-checking. This transformation is also evident in the shift of claim prioritization from focusing solely on checking political claims to addressing online misinformation by considering a broader range of checkworthy factors (described in Section 2.1). Thus, claim prioritization has become a central challenge for fact-checkers. In this section, by illustrating the historical changes in claim prioritization and its differences across various fact-checking entities, we discuss how our study insights better inform the development of advanced tools in order to adapt to this evolving practice.
Traditionally, fact-checking was an ad-hoc process embedded within news production prior to publication, where editors worked closely with authors to verify the accuracy of arguments (Smith, 2003). Ideally, every claim in ready-to-publish news was considered important to fact-check as part of internal journalistic work. This helps maintain the credibility and reputation of the media outlet. As fact-checking has evolved into a post-hoc activity focused on assessing the accuracy of existing public statements, claim prioritization has become a distinct step. In this step, journalists start assessing the newsworthiness of various claims and only select some of them that are worth checking. While this step is recognized as important, it is not generally viewed as particularly challenging. For example, as highlighted in the ethnographic work of Graves (2016), post-hoc fact-checking was largely carried out by traditional newsrooms and media outlets that primarily focused on political claims. Journalists in these organizations often focus on statements made by politicians, other journalists, and public figures. This emphasis remains evident, as noted by some participants (P5, 12) and other related work (Mena, 2019; Micallef et al., 2022).
@@
Claim prioritization tool has the potential to accelerate the collaborative effort to meet the fact-checking needs of different stakeholders. For example, by using the tool to prioritize claims based on organizational checkworthiness, fact-checkers across these organizations can enhance their capacity to address a wider range of claims. Additionally, the tool might help address existing ethical considerations, particularly when fact-checkers navigate competing objectives in selecting claims. According to our participants, these competing objectives often stem from their partnerships with social media platforms and regional differences. We discuss these in the following sections.